OpenStudy (anonymous):

What are the x-intercept(s) of the graph of the equation y - x2 = -4x - 5? (1 points) (1, 0) and (-5, 0) (1, 5) (-1, 0) and (-5, 0) (-1, 0) and (5, 0)

OpenStudy (anonymous):

to find the x-int you have to set y = 0 so now you have 0 = x^2 - 4x -5 now just solve the regular quadratic equation! and you will have your answers!

OpenStudy (anonymous):

ohhhhh...... the x^2 on the other side of the equal sign got me confused but I got it haha thanks^.^

OpenStudy (anonymous):

no problem :)

OpenStudy (anonymous):

can u help me with another problem?

OpenStudy (anonymous):

i will try my best :)

OpenStudy (anonymous):

thanks(:Solve: 0 = 6x^2 - 10x - 4. (1 points)

OpenStudy (anonymous):

x=-1/3, x=-2 x=-1/3 x=2 x=1/3 x=-2 x=1/3 x=2

OpenStudy (anonymous):

 i got you started :)